a crane a force ,F to lift a 200kg load up into the air . the load is accelerating upwards at o.5 m.s-2 . calculate the value of the applied force , F
1
Expert's answer
2020-11-12T09:24:55-0500

Answer

By second law of motion

"F-mg=ma\\F=m(g+a)"

Mass(m) =200kg

Acceleration (a) =0.5m/s^2

So

"F=200\times (9.8+0.5)\\ F=2060N"